Patel and Sons Inc. uses a standard cost system to apply factory overhead costs to units produced. Practical capacity for the plant is defined as 52,200 machine hours per year, which represents 26,100 units of output. Annual budgeted fixed factory overhead costs are $261,000 and the budgeted variable factory overhead cost rate is $2.70 per unit. Factory overhead costs are applied on the basis of standard machine hours allowed for units produced. Budgeted and actual output for the year was 19,700 units, which took 41,200 machine hours. Actual fixed factory overhead costs for the year amounted to $253,800 while the actual variable overhead cost per unit was $2.60.

Based on the information provided above, calculate the following factory overhead variances for the year. Indicate whether each variance is favorable (F) or unfavorable (U). (Do not round intermediate calculations. Round your final answers to nearest whole dollar amount.)

a. Total overhead Variance
b. Total flexible-budget Variance
c. Production Volume Variance
Answer #1

Standard time for producing one unit of output = 52,200 / 26,100 = 2 hours per unit of output

Budgeted Fixed Overhead = $2,61,000    

Budgeted overhead absorption rate per hour = $2,61,000 / 52,200 hours = $5 per hour.

Budgeted units produced during the year was 19700 units. Standard hours for 19,700 units = 2 x 19,700 = 39,400 hours.

a) Total Fixed Overhead Variance = standard hours allowed for actual production x standard overhead absorption rate per hour - Actual Fixed overhead cost incurred

= 39,400 hrs. x $5 per hour - $2,53,800 = $56,800 (U).

Standard Variable Overhead rate = $2.70 per unit.

Variable Overhead Variance = Actual output x Standard rate per unit - Actual variable overhead

= 19,700 units x $ 2.70 per unit – 19,700 units x $2.60 per unit = 1,970 (F).

Total Overhead Variance = Total Fixed Overhead Variance + Total Variable Overhead Variance = $56,800 (U) + $1970 (F) = $54,830 (U).

b) When flexible budget will be prepared for actual production of 19700 units, the fixed expense will remain same as the budget i.e. $2,61,000.

Flexible overhead budget total for actual output of 19,700 units = $2,61,000 + 19,700 x $2.70

= $3,14,190.

Actual overhead expense = $2,53,800 + 19,700 x $2.60 = $3,05,020.

Total Flexible budget variance = $3,14,190 - $3,05,020 = $9,170 (F).

c) Production Volume Variance

= (Budgeted Hours - Standard Hours required for actual production) x Standard overhead rate per hour = (52,200 hrs – 19,700 units x 2 hrs /unit) x $5/hr = $64,000